OpenStudy (howard-wolowitz):

Microsoft offers its employees one of three pension splans, A, B, or C. Past history shows that 70% of employees choose plan A, 20% choose plan B, and 10% choose plan C. Also, 60% of those who choose plan A are married, 20% of those who choose plan B are married, and 60% of those who choose plan C are married. If a married employee is selected at random, what is the probability the employee is in plan A?

OpenStudy (howard-wolowitz):

i think the answer is 0.81

hartnn (hartnn):

Baye's theorem. \(\dfrac{0.7\times 0.6 }{0.7\times0.6+ 0.2\times 0.2+0.1\times 0.6}\)

hartnn (hartnn):

you're correct :)

OpenStudy (howard-wolowitz):

thakn you for helping

hartnn (hartnn):

welcome ^_^